A nutrition plan counts grams of fat, carbohydrates, and fiber. The grams of carbohydrates are to be 15 more than twice the grams of fat. The grams of fiber are to be three less than the grams of fat. If the grams of carbohydrate, fat, and fiber must total 48 grams for a dinner meal, how many grams of each would be in the meal?
Answer by richwmiller(17219)

t for fat
r for fiber
c=15+2t
t-3=r
r+t+c=48
c = 33, r = 6, t = 9